OpenStudy (anonymous):

The height in feet reached by a baseball tossed upward at a speet of 48 ft/second from the ground is given by h(t)=-16t^2+48t, where t represents time in seconds after the ball is tossed. at what time does the baseball reach 24ft? use the quadratic formula... ok I know this can be solve using the vertex...right? how do you solve it?

OpenStudy (anonymous):

The LHS of the equation is your height. If you want to calculate t for a given height, write: 24 = -16t^2 + 48t. And solve the quadratic equation using the quadratic formula. The negative solution does not make physical sense, so pick the positive one.

OpenStudy (campbell_st):

equate the information then 24 = -16t^2 + 48t then 16t^2 - 48t + 24 = 0 or 8(2t^2 - 6t + 3) = 0 you need to solve for t there will be 2 time values as the path is parabolic... there will be a time going up and going down you need the general quadratic formula to show \[t = (6 \pm2\sqrt{3})4\]

OpenStudy (campbell_st):

oops \[t = (6 \pm 2\sqrt{3})/4\]

OpenStudy (anonymous):

Just a remark: when I said the negative solution does not make physical sense what I meant is that if you end up with a negative value for time, that doesn't make that much of a sense. In this problem, both must be positive because you pass at h = 24ft two times: going up and going down.

OpenStudy (anonymous):

what if do this...would it be wrong?

OpenStudy (campbell_st):

thats the axis if symmetry of the curve and the vertex of the curve lies on that line

OpenStudy (campbell_st):

but the vertex won't help in solving the problem other than letting you know there is a value for t either side of t = 3/2

OpenStudy (anonymous):

oh ok i get it it would just give me the height not the time..